# technical note on calibrating vision-language models under covariate shift

### v1

@@ 3 Experiments and Results
4 Conclusion
@@
This work introduces C3SC a unified framework to address two critical challenges covariate shift and confidence misalignment in low-shot setting vision tasks with foundation model. Integration of Fisher information and CMP, significantly improve both robustness and calibration across various vision and domain generalization datasets. Experimental results demonstrate that C3SC outperforms existing methods in both accuracy and calibration consistently across datasets (upto $3.5\%$) on challenging covariate shift datasets) and calibration (reducing Expected Calibration Error by up to $5.82\%$). These findings highlight C3SC as a promising solution for reliable real-world applications, particularly in low-shot settings where distribution shifts and overconfidence are prevalent.
Author Contributions
If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors.
@@ Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
Appendix A Appendix
A.1 Theoretical Background
Proposition A.1.
@@ A.1 Theoretical Background
$\log p(X;\theta)\approx\log p(X;\theta_{q})+(\theta-\theta_{q})^{T}\nabla_{\theta}\log p(X;\theta_{q})+\frac{1}{2}(\theta-\theta_{q})^{T}H(\theta-\theta_{q}),$
@@
Thus, Fisher information provide a measure of divergence between $q(X)$ and $p(X;\theta)$, capturing covariate shift. ∎
Proposition A.2.
@@
Thus, CMP acts as confidence misalignment penalty, discouraging overconfident yet incorrect predictions. ∎
Proposition A.3.
@@
$\mathcal{L}_{\text{$C3SC$}}(x,y;\theta)=\mathcal{L}_{c}+\lambda_{1}I(\theta)+\lambda_{2}\text{CMP}$
@@
Thus, minimizing $\mathcal{L}_{\text{$C3SC$}}$ ensures well-calibrated robust prediction under covariate shift. ∎
Appendix B Experiments
Datasets. To check the effectiveness of $C3SC$, we used 19 vision dataset including natural distribution datasets, domain adaptation datasets and vision classification datasets used for evaluation by low-shot prompt learning methods. These datasets are: Caltech101 Fei-Fei et al. [2004], Imagenet Deng et al. [2009], EuroSAT Helber et al. [2019], StanfordCars Krause et al. [2013], FGVCAircraft Maji et al. [2013], OxfordPets Parkhi et al. [2012], SUN397 Xiao et al. [2010], Food101 Bossard et al. [2014], Flowers102 Nilsback and Zisserman [2008], DTD Cimpoi et al. [2014], UCF101 Soomro [2012], PACS Li et al. [2017], VLCS Fang et al. [2013], Office-Home Venkataraman et al. [2016] and five domain adaptation datasets such as Imagent-1k and its variants like ImageNet A (Art) Hendrycks et al. [2021a], ImageNet-V2 (V2) Recht et al. [2019], ImageNet R (Real) Hendrycks et al. [2021b] and ImageNet S (Sketch) Wang et al. [2019].
Baselines. We consider CLIP and prompt learning method CoOpZhou et al. [2022] without any penalty integration as baseline in comparison to our method.
Evaluation metrics. We used accuracy (Acc) as evaluation metric for robustness of $C3SC$ under covariate shift and vision datasets used by prompt learning approaches, while Expected Calibration Error (ECE) Guo et al. [2017] is used with same number of datasets for evaluating the calibration performance of $C3SC$.
Implementation details. To check the robustness and calibration performance of $C3SC$, we conduct extensive experiments on various datasets. We used pre-trained CLIP Radford et al. [2021] with ViT-B/16 Dosovitskiy [2020] as backbone in prompt learning method. In covariate shift experiment we used CLIP with ViT-B/16 and ResNet-50 backbone He et al. [2016] for the last four variants of ImageNet. For first four datasets we follow the same protocol of Li et al. [2017], where model is trained on another domain and evaluated on another test domain.
Table 3: $C3SC$ accuracy results on covariate shift vision datasets with and without FIM penalty. The $\Delta$ row shows the percentage increase/decrease in accuracy. $\uparrow$ shows improvement in accuracy while the $\downarrow$ shows decrease in performance.
CLIP 96.1 80.4 81.4 54.1 60.8 46.2 47.8 73.9
CoOp 96.5 82.1 82.5 58.8 64.2 47.9 49.7 75.2
CoOp+FIM 98.0 85.6 86.0 60.0 65.5 48.8 50.5 76.8
Table 4: $C3SC$ ECE results on covariate shift vision datasets with and without CMP penalty. The $\Delta$ row shows the percentage increase/decrease in calibration. $\downarrow$ shows improvement in calibration while $\uparrow$ the shows calibration performance decrease.
ECE Method PACS Office-Home VLCS DomainNet ImageNet-V2 ImageNet-S ImageNet-A ImageNet-R
CLIP 2.18 2.77 2.89 3.69 2.44 4.88 8.34 3.51
CoOp 2.02 2.92 3.01 3.29 4.19 8.40 15.34 3.12
CoOp+CMP 1.91 2.75 2.84 3.15 4.05 8.18 15.00 2.95
Table 5: Accuracy three datasets for tuning $\boldsymbol{\lambda_{1}}$ within range ($0.0-1.0$) while keeping $\boldsymbol{\lambda_{2}}$ value fixed $0.1$.

### v2

@@ 3 Method
$\mathcal{L}_{\text{CalShift}}(x,y;\theta)=\mathcal{L}_{c}+\lambda_{1}I(\theta)+\lambda_{2}\text{CMP}$ (3)
More detail on the theoretical background about $I(\theta)$ and CMP can be found in Appendix A.
4 Experiments and Results
4.1 Experiments
Datasets. To check the effectiveness of CalShift, we used 19 vision dataset including natural distribution datasets, domain adaptation datasets and vision classification datasets used for evaluation by low-shot prompt learning methods. These datasets are: Caltech101 [12], Imagenet [9], EuroSAT [18], StanfordCars [29], FGVCAircraft [33], OxfordPets [43], SUN397 [62], Food101 [5], Flowers102 [38], DTD [6], UCF101 [49], PACS [32], VLCS [11], Office-Home [54] and five domain adaptation datasets such as Imagent-1k and its variants like ImageNet A (Art) [22], ImageNet-V2 (V2) [46], ImageNet R (Real) [21] and ImageNet S (Sketch) [56].
Baselines. We consider CLIP and prompt learning method CoOp[67] without any penalty integration as baseline in comparison to our method.
Evaluation metrics. We used accuracy (Acc) as evaluation metric for robustness of CalShift under covariate shift and vision datasets used by prompt learning approaches, while Expected Calibration Error (ECE) [15] is used with same number of datasets for evaluating the calibration performance of CalShift.
Implementation details. To check the robustness and calibration performance of CalShift, we conduct extensive experiments on various datasets. We used pre-trained CLIP [45] with ViT-B/16 [10] as backbone in prompt learning method. In covariate shift experiment we used CLIP with ViT-B/16 and ResNet-50 backbone [17] for the last four variants of ImageNet. For first four datasets we follow the same protocol of [32], where model is trained on another domain and evaluated on another test domain.
The experimental detail, datasets, and evaluation metrics used in the paper can be found in appendix 4.1. here,
we discuss the robustness and confidence calibration results of our method CalShift.
4.2 Results
Calshift generalization performance across shots. Table 1 upper part shows the generalization performance of CalShift across incremental few-shot settings. It is shown in Table 1 that the integration of FIM with CoOp consistently outperforms CLIP and vanilla CoOp across all shot settings. CalShift achieves a maximum accuracy improvement of $6.8\%$ over CoOp in the zero-shot setting and $7.2\%$ in the 1-shot setting. CalShift shows a $5.9\%$ average accuracy improvement over CoOp, demonstrating that CalShift regularization effectively mitigates overfitting in data-scarce regimes.
The $\Delta\%$ decrease monotonically from $7.2\%$ ($1$-shot) to $5.1\%$ ($16$-shot) settings, which shows the effectiveness of CalShift in extremely low-shot data regime. This provides empirical strength to FIM’s role in stabilizing gradient updates when labeled examples are scarce. Furthermore, in a zero-shot setting with no training data, integrating FIM with CoOp surpasses CLIP, demonstrating FIM’s superior ability to retain knowledge compared to CoOp alone. This suggests that FIM acts as a prior-preserving regularizer, preventing CLIP from catastrophic forgetting in zero-shot scenarios.
Calshift calibration performance across shots. CalShift’s calibration performance results are shown in the lower half of Table 1. Integrating CMP with CoOp consistently outperforms vanilla CoOp across all few-shot settings, achieving up to a $12.5\%$ improvement in the $4$-shot setting and $10.2\%$ in the $16$-shot scenario. In the zero-shot scenario, CLIP’s calibration performance surpasses CoOp, indicating that prompt tuning can degrade CLIP’s calibration. Integrating CMP into CoOp mitigates this issue, improving calibration by $8.93\%$ in the zero-shot setting.
Integrating CMP into CoOp improves $\Delta\%$ by a maximum of $8.97\%$ in the $1$-shot setting and $10.2$% in the 16-shot setting, demonstrating CMP’s ability to better leverage additional data for confidence calibration. Table 1 demonstrates that integrating CMP into CoOp outperforms both vanilla CoOp and CLIP beyond the 4-shot setting, suggesting that CMP enhances calibration adaptation. This suggests that CMP tackles CoOp’s two key issues by penalizing overconfident logits: (1) overfitting in few-shot scenarios and (2) gradient misalignment . CMP addresses overfitting by enforcing decision boundary constraints, preventing overconfidence on limited data. Meanwhile, it mitigates misalignment by aligning logit gradients with true class margins, reducing bias in probability estimates.
Table 1 shows that FIM integration significantly enhances generalization, particularly in low-shot regimes, while CMP integration improves calibration in these settings. These results establish CalShift as a versatile regularizer, ensuring robust and trustworthy adaptation in vision-language models.
Table 3: The upper part shows CalShift accuracy results on covariate shift vision datasets with and without FIM penalty. The $\Delta$ row shows the percentage increase/decrease in accuracy. $\uparrow$ shows improvement in accuracy while the $\downarrow$ shows decrease in performance. The lower part shows CalShift ECE results on covariate shift vision datasets with and without CMP penalty. The $\Delta$ row shows the percentage increase/decrease in calibration. $\downarrow$ shows improvement in calibration while $\uparrow$ shows calibration performance decrease.
CLIP 96.1 80.4 81.4 54.1 60.8 46.2 47.8 73.9
CoOp 96.5 82.1 82.5 58.8 64.2 47.9 49.7 75.2
CoOp+FIM 98.0 85.6 86.0 60.0 65.5 48.8 50.5 76.8
ECE
CLIP 2.18 2.77 2.89 3.69 2.44 4.88 8.34 3.51
CoOp 2.02 2.92 3.01 3.29 4.19 8.40 15.34 3.12
CoOp+CMP 1.91 2.75 2.84 3.15 4.05 8.18 15.00 2.95
CalShift robustness in vision tasks. The results in Table 2 demonstrates the effectiveness of FIM penalty integration into CoOP as compared to baselines. The $\Delta$ row in Table 2 shows improvement across most datasets. The notable improvement in accuracy include Imagenet ($6.8\%$), Flowers102102102102 ($8.0\%$), and UCF101101101101 ($7.3\%$), indicates significant improvement, from action recognition to fine-grained classification. In scene and texture-based datasets CalShift shows robustness performance improvement with SUN397397397397 ($6.0\%$) and DTD ($4.9\%$). For Food101101101101 ($1.7\%$), Caltech101101101101 ($0.2\%$), and OxfordPets ($0.7\%$) improvement in accuracy. CalShift improves in accuracy by $3.2\%$ on average across all datasets.
CalShift calibration in vision tasks. The CMP penalty-based calibration performance results across eleven vision datasets are shown in Table 2. In Table 2, it is shown that CMP integration improves in calibration for most datasets with an average of $5.70\%$ reduction in calibration.
Smaller values are better in the calibration case. Penalty-based CoOp shows $9.85\%$ calibration improvement at maximum on Food101, mageNet ($8.93\%$), and OxfordPets ($7.19\%$), which demonstrates CMP alignment in model calibration in these settings. On other datasets also it shows better performance such as, UCF101101101101 ($4.55\%$), Caltech101 ($4.94\%$), StanfordCars ($6.51\%$), and FGVCAircraft ($4.94\%$).
Comparing penalty-bases CoOp to CLIP, penalty-based enhancement provides significant improvement across almost all datasets except StanfordCars and EuroSAT. CalShift shows UCF101101101101 ($14.4\%$), Flowers102102102102 ($13.4\%$), and ImageNet ($12.5\%$) improvement at maximum in comparison to CLIP.
The integration of FIM penalty generally shows improvement in accuracy in low-shot vision tasks in particular when combined with prompt learning methods.
CalShift robustness in covariate shift. Table 3 shows the robustness performance of CalShift on eight domain generalization datasets under covariate shift. The $\Delta$ of Table 3 shows the penalty-based CoOp outperform CoOp across all datasets, showing significant improvement in robustness to covariate shift. CalShift shows maxmimum robustness performance in Office-Home ($3.5\%$) and VLCS ($3.5\%$) demonstrating FIM integration enhances adaptation across domain adaptation tasks. For all other datasets such as, DomainNet ($1.2\%$), ImageNet-V2 ($1.3\%$), ImageNet-S ($0.9\%$), ImageNet-A ($0.8\%$), and ImageNet-R ($1.6\%$), CalShift shows slight better improvement, suggesting better robustness in challenging real-world distribution shift.
While comparing penalty-based CoOp to CLIP, we observe consistent improvement in accuracy across all domain generalization datasets like, VLCS ($4.6\%$) , Office-Home ($5.2\%$) and PACS ($1.9\%$). CalShift demonstrates better robustness on more challenging variants of ImageNet such as, ImageNet-V2 ($1.3\%$) , ImageNet-S ($2.6\%$), ImageNet-A ($2.7\%$), and ImageNet-R ($2.9\%$).
CalShift calibration in covariate shift. Table 3 presents, calibration performance of penalty integration into CoOp for covariate shift vision datasets. It is demonstrated, that CalShift improves in calibration across all datasets with an average of $2.21\%$ at minimum and $5.82\%$ at maximum. CalShift shows calibration improvement on Office-Home ($5.82\%$), PACS ($5.45\%$), VLCS ($5.64\%$), ImageNet-R ($5.45$%) DomainNet ($4.26\%$) and ImageNet-V2 ($3.34\%$), while smaller but consistent improvements are seen in ImageNet-S ($2.62\%$) and ImageNet-A ($2.21\%$), which indicates the CMP effectiveness in confidence alignment on these datasets.
CLIP shows better calibration as compared to CoOP by reporting lower ECE such as, PACS ($2.18$ vs. $2.02$), VLCS ($2.89$ vs. $3.01$), and ImageNet-R ($3.51$ vs. $3.12$). However, when penalty is integrated into CoOp, the calibration reduces significantly making it more comparable to CLIP.
Ablation of loss components. We evaluate the performance of our proposed architecture by conducting ablation studies to understand the relative contribution of each component, Fisher information and CMP. We ablate the model by keeping the Fisher information $I(\theta)$ hyperparameter $\lambda_{1}=0$ in Equation 3 and then perform the experiment. Similarly, we set the CMP hyperparameter $\lambda_{2}=0$ and conduct another experiment. The results are reported in Table 4 and Table 5, respectively.
Effect of Fisher Information. Table 4 evaluates the isolated contributions of CMP and FIM across 11 vision datasets. The upper part of the table presents CalShift’s generalization performance with $\lambda_{1}=0$ and $\lambda_{2}=0.4$. CalShift demonstrates generalization efficacy, achieving a maximum improvement of $3.3\%$ on Flowers102, $1.9\%$ on SUN397, and $2.6\%$ on ImageNet. However, it shows only marginal accuracy gains on OxfordPets and Caltech101, suggesting that CMP’s effect diminishes when CLIP features are already sufficient.
The lower part of Table 4 presents CalShift’s accuracy across 11 vision datasets with $\lambda_{1}=0.4$ (FIM-only) and $\lambda_{2}=0$ (no-CMP). CalShift achieves a maximum accuracy improvement of $5.2\%$ on FGVCAircraft, $7.4\%$ on EuroSAT, and $4.9\%$ on Flowers102102102102.
Effect of misalignment penalty. Table 5 shows isolated calibration effect of CalShift across 11 vision datasets. The upper section of the Table shows ECE results with $\lambda_{1}=0$ (no FIM ) and $\lambda_{2}=0.4$ (CMP only). CalShift improves in ECE by 2.7% on Flowers102 suggesting Calshift align well with balanced label distributions, However, limited gains on CLIP-compatible datasets 1.96% on OxfordPets and 1.3% on EuroSAT.
The lower section of the Table 5 with $\lambda_{1}=0.4$ (FIM-only) and $\lambda_{2}=0$ (no-CMP). It is shown in Table that CalShift consistently enhaces calibration, reducing average ECE by $1.90\%$. Notable improvements occur on fine-grained datasets like StanfordCars -$1.78\%$ and class-balanced benchmarks such as Flowers102: -$4.05\%$, demonstrating FIM ability to stabilize confidence estimates by penalizing overparameterized gradients.
Effect of $\lambda_{1}$ tuning on CalShift performance. Table 6 given in appendix B shows accuracy performance on four datasets, including Flowers102102102102, Food101101101101, UCF101101101101, and DTD, for different values of $\lambda_{1}$ within the range (0.0 to 1.0), while $\lambda_{2}$ remains fixed at 00. $\lambda_{1}=0$ serves as the baseline with no Fisher penalty. The results in the table indicate that CalShift achieves the highest performance at $\lambda_{1}=0.4$, with accuracies of $98.8\%$ on Food101101101101, $84.3\%$ on UCF101101101101, and $55.1\%$ on DTD, whereas performance declines for all other values.
Effect of $\lambda_{2}$ tuning on CalShift performance. In appendix Table 7 shows the calibration results of CalShift when tuning $\lambda_{2}$ within range (0.0 to 1.0) while $\lambda_{1}$ value fixed to 0. It is show in Table that with $\lambda_{2}=0.4$ CalShift achieve better calibration results across four dataset. The Calshift calibration performance drops on other values of $\lambda_{2}$, while value $0.4$ balance between over-regularization and under-regularization.
Our method is effective in confidence alignment as it shows better ECE consistently on almost all datasets. These results favors CalShift calibration reliability under covariate shift, making it more trustworthy in real-world applications.
5 Conclusions
@@
This work introduces CalShift, a unified framework designed to address two critical challenges in low-shot vision tasks using foundation models: covariate shift and confidence misalignment. By integrating Fisher Information and Confidence Misalignment Penalty (CMP), CalShift significantly enhances both robustness and calibration across various vision and domain generalization benchmarks. Experimental results show that CalShift (Confidence-Calibrated Covariate Shift Correction), built within the CalShift framework, consistently outperforms existing methods in terms of accuracy and calibration. Specifically, it achieves up to a 3.5% gain in accuracy and reduces Expected Calibration Error (ECE) by up to 5.82% on challenging covariate shift datasets. These findings underscore CalShift’s promise as a reliable solution for real-world deployment, especially in low-shot scenarios where distribution shifts and overconfidence are prevalent. Future directions include extending the framework to continual learning, online adaptation, and test-time adaptation, where the model must remain calibrated under evolving distributions.
CalShift is well-suited for scenarios like test-time adaptation where model adapts to unseen (from a ststistics point of view, shifted) data distribution at inference time. The integration of Fisher information and CMP provides robust adaptation to distribution shift without need of additional training data. This makes it an effective solution particularly in those real-world applications where test-time condition vary from those during training.
The issue of confidence misalignment is not just a technical challenge but a critical challenge in AI alignment, leading to unsafe and unreliable decisions in safety-critical systems such as medical diagnosis and autonomous vehicles. Our method CalShift ensures that model predictions is aligned with human expectation.
@@ Acknowledgments
The lead author experimented at NCBC-Video Surveillance Lab at FAST-NUCES Karachi, led by Prof. Muhammad Atif Tahir.
Appendix A Theoretical Background
Proposition A.1
@@ Appendix A Theoretical Background
$\log p(X;\theta)\approx\log p(X;\theta_{q})+(\theta-\theta_{q})^{T}\nabla_{\theta}\log p(X;\theta_{q})\\ +\frac{1}{2}(\theta-\theta_{q})^{T}H(\theta-\theta_{q})$
@@
Thus, Fisher information provide a measure of divergence between $q(X)$ and $p(X;\theta)$, capturing covariate shift.
Proposition A.2
@@
Thus, CMP acts as confidence misalignment penalty, discouraging overconfident yet incorrect predictions.
Proposition A.3
@@
$\mathcal{L}_{\text{CalShift}}(x,y;\theta)=\mathcal{L}_{c}+\lambda_{1}I(\theta)+\lambda_{2}\text{CMP}$
@@
Thus, minimizing $\mathcal{L}_{\text{CalShift}}$ ensures well-calibrated robust prediction under covariate shift.
Appendix B Ablation study
Table 4: The upper part of the table shows CalShift accuracy results while keeping Fisher Information penalty $I(\theta)$ ($\lambda_{1}=0$) and CMP penalty ($\lambda_{2}=0.4$). The $\Delta$ row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy compared to CoOp. The lower part of the table CalShift accuracy results while keeping CMP penalty ($\lambda_{2}=0$) and FIM penalty $\lambda_{1}=0.4$. The $\Delta$ row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy compared to CoOp.
CLIP 69.9 90.1 96.8 91.2 72.1 72.4 63.3 27.2 69.4 53.3 56.5 69.3
CoOp 78.6 97.0 98.6 98.2 79.2 79.5 59.2 25.2 63.0 52.5 53.8 71.2
CoOp + CMP ($\lambda_{1}=0$) 81.5 97.5 98.6 98.4 81.8 81.6 57.8 26.0 64.2 53.6 51.2 71.9
$\Delta$ 3.7 ↑ 0.5 ↑ 0.0 → 0.2 ↑ 3.3 ↑ 2.6 ↑ 2.4 ↓ 3.2 ↑ 1.9 ↑ 2.1 ↑ 4.8 ↓ 1.0 ↑
CoOp 78.6 97.0 98.6 98.2 79.2 79.5 59.2 25.2 63.0 52.5 53.8 71.2
CoOp + FIM ($\lambda_{2}=0$) 82.3 97.9 98.5 98.7 83.1 82.9 57.2 26.5 65.8 54.5 49.8 72.6
$\Delta$ 4.7 ↑ 0.9 ↑ 0.1 ↓ 0.5 ↑ 4.9 ↑ 4.3 ↑ 3.4 ↓ 5.2 ↑ 4.4 ↑ 3.8 ↑ 7.4 ↓ 2.0 ↑
Table 5: The upper half of the table shows CalShift ECE results on vision datasets with and without FIM penalty $\lambda_{1}=0$ and CMP penalty ($\lambda_{2}=0.4$). The $\Delta$ row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in calibration error. CalShift ECE results on vision datasets with and without CMP penalty ($\lambda_{2}=0$) and FIM penalty ($\lambda_{1}=0.4$) . The $\Delta$ row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in calibration error.
CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16
CoOp + CMP ($\lambda_{1}=0$) 3.18 3.42 3.32 3.12 2.88 3.48 3.50 3.32 3.08 3.12 3.12 3.24
$\Delta$ 3.25$\uparrow$ 2.09$\uparrow$ 2.47$\uparrow$ 1.96$\uparrow$ 2.70$\downarrow$ 3.57$\uparrow$ 3.55$\uparrow$ 2.47$\uparrow$ 1.99$\uparrow$ 1.96$\uparrow$ 1.30$\uparrow$ 2.53$\uparrow$
CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16
CoOp + FIM ($\lambda_{2}=0$) 3.05 3.30 3.20 3.00 2.84 3.30 3.32 3.18 2.96 3.00 3.02 3.10
$\Delta$ 0.97$\downarrow$ 1.49$\downarrow$ 1.23$\downarrow$ 1.96$\downarrow$ 4.05$\downarrow$ 1.79$\downarrow$ 1.78$\downarrow$ 1.85$\downarrow$ 1.99$\downarrow$ 1.96$\downarrow$ 1.94$\downarrow$ 1.90$\downarrow$
Table 6: Accuracy three datasets for tuning $\boldsymbol{\lambda_{1}}$ within range ($0.0-1.0$) while keeping $\boldsymbol{\lambda_{2}}$ value fixed 00.
@@ Appendix B Ablation study
Table 7: Expected Calibration Error (ECE) for tuning $\boldsymbol{\lambda_{2}}$ in range ($0.0-1.0$) with $\boldsymbol{\lambda_{1}}=0$.
$\boldsymbol{\lambda_{2}}$ Datasets
Flowers102 Food101 UCF101 DTD
0.0 6.27 5.69 5.21 4.13
0.2 4.61 4.36 4.04 3.66
0.4 3.16 3.02 2.94 2.92
0.6 3.56 3.54 3.24 3.15
0.8 4.26 4.18 3.74 3.39
1.0 5.22 5.17 4.48 3.89